OpenStudy (anonymous):

the square root of x+4 = x-8? I'm doing a summer math packet and this is one problem that really stumped me...

hartnn (hartnn):

to get rid of square root,square both sides because \[(\sqrt{a})^{2}=a\]

OpenStudy (anonymous):

can you explain that a bit more?

OpenStudy (anonymous):

since x+4 is already squared, would i also square x-8?

hartnn (hartnn):

your question is \[\sqrt{x+4}=x+8\] isn't it? squaring both sides will give \[x+4=(x+8)^{2}\]

OpenStudy (anonymous):

right...

OpenStudy (anonymous):

then, would i say: x+4 = (x-8)(x-8)?

hartnn (hartnn):

yes,go on and multiply (x-8)(x-8)

OpenStudy (anonymous):

using the FOIL method, right?

hartnn (hartnn):

u work it out and show the result,i'll tell u whether its correct...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

x+4=x squared - 8x - 8x + 64 then... x+4=x squared -16x + 64

OpenStudy (anonymous):

\[x ^{2} -16x +64\]

OpenStudy (anonymous):

right. so, what do i do next?

hartnn (hartnn):

yes thats correct for (x-8)(x-8) now equate it to x+4 combine like terms and form a quadratic equation...try

OpenStudy (anonymous):

what do you mean by "equate it to"?

hartnn (hartnn):

\[x+4=x ^{2}-16x+64\] i meant this

hartnn (hartnn):

now can u combine terms with x and constants?

OpenStudy (anonymous):

answer is x=12 i got it only by checking.yet one another value must satisfy the same relation because is quadratic equation

OpenStudy (anonymous):

oh, yeah, i already got that when i solved: (x-8)(x-8)

OpenStudy (anonymous):

combine terms? do you mean combining x with -16x and 4 with 64?

OpenStudy (anonymous):

i got x squared - 15x + 68

OpenStudy (anonymous):